# An open-source heuristic to reboot 2D nesting research

### v1

@@ 8.3 Balancing exploration and exploitation
2:move_items $\triangleright$ Alg. 5
[…]
6:restore to $s^{*}$
@@ 9 Solving the 2D irregular strip packing problem
[…]
@@ 9.1 Exploration phase
[…]
4:shrink strip by $\texttt{R}_{x}$
[…]
@@
The separation procedure (Algorithm 9) attempts to convert this infeasible solution into a feasible one. If successful, the returned solution $s$ replaces the incumbent solution $s^{*}$, after which the strip is once again shrunk. If unsuccessful, $s$ represents an infeasible local optimum which is added to a pool of solutions $\mathcal{S}$. For the next separation attempt, a random solution $\hat{s}$ is chosen from $\mathcal{S}$, with those closer to feasibility (Equation 8) having a greater chance of being selected. A new starting solution $s^{\prime}$ is then created by swapping two large items from $\hat{s}$. This ensures $s^{\prime}$ inherits many high-quality placements from $\hat{s}$ while, at the same time, being sufficiently disrupted so it can escape from the local optimum. This procedure continues until time limit $\texttt{TL}_{x}$ is reached, at which point the exploration phase returns the best feasible solution found $s^{*}$.
@@ 9.2 Compression phase
$\texttt{TL}_{c}$not reached
[…]
3:restore to $s^{*}$ and shrink strip by $r$
[…]
@@
Shrink ratio $r$ is much smaller than during the exploration phase and decays linearly from $\texttt{R}_{c}^{s}$ to $\texttt{R}_{c}^{e}$ based on the phase’s remaining time limit $\texttt{TL}_{c}$. This means that the separate procedure only needs to resolve (progressively) less severe collisions, which it is often able to do without making major changes to the configuration of items. Because the procedure always restores the best feasible solution and then shrinks the strip, the search process’ freedom is restricted. The compression phase is therefore focused on finding out how much further the strip can be shrunk without deviating too far from the configuration of items provided by the exploration phase.
@@ 10 Implementation
The implementation of our 2D nesting algorithm is written in Rust and called sparrow. It is publicly available at: https://github.com/JeroenGar/sparrow. The codebase closely follows the structure of this paper, making it straightforward to understand how sparrow relates to the algorithms introduced in this paper. Throughout this project, we have strived for maximum computational performance, robustness and leanness of both the algorithm and its implementation.
@@
So far, this limitation concerning exact fits only surfaces in artificial academic instances featuring simple shapes and vertices with integer coordinates. Conditions which, generally, do not occur in real-world applications.
@@ 11.1 Experimental setup
We used a combination of manual tuning and hyperparameter optimization (Bergstra et al., 2013) to identify a suitable set of parameters for the algorithm. These parameters were optimized for the maximum expected performance across all instances from the academic benchmark set when subject to a 20-minute time limit. This single set of parameters was used for every instance in our experiments.
@@ 11.2 Performance analysis
To evaluate the performance of sparrow we will use academic benchmark instances from the ESICUP website [^3] https://www.euro-online.org/websites/esicup/data-sets. We refer to Table 3 in Sato et al. (2019) for an overview of these instances and their properties.
@@
The expected quality $\mathbb{E}(\rho)$ and interquartile range of sparrow’s solutions in comparison to the academic SotA algorithms for the 2DISPP and one open-source implementation are presented in Table 2. The distributions of the solutions from Table 2 are visualized as violin plots in Figure 9. One important thing to note is that the width of the strip for some of the most simple instances – SHAPES0, SHAPES1, SHAPES2, FU, JAKOBS1 and JAKOBS2 – was inflated by 0.01%. This minuscule inflation should not affect which configurations of items are possible, but suffices to avoid the issue concerning exact fits (described in Section 10) which would otherwise occur in these artificial instances. For academic purposes, the best solution produced by sparrow for each instance, alongside the previous best-known solution, is presented in Table 3 and shown in Figure 10.
@@
Table 2: The expected solution quality $\mathbb{E}(\rho)$ on academic benchmarks obtained by sparrow, four algorithms from the literature: ROMA - Sato et al. (2019), GCS - Elkeran (2013), FLD - Wang et al. (2017), ELS - Leung et al. (2012) and one open-source implementation: PS - Fontan (2023). For sparrow, quartiles (Q1, Q3) and the margin of error ($\pm$) on $\mathbb{E}$, defined as the half-width of the 95% confidence interval computed via non-parametric bootstrapping, are reported. The reported values for sparrow result from 100 independent runs, each with a time limit of 20 minutes. Values for ROMA, GCS, FLD and ELS were taken from their respective publications. PS is a deterministic algorithm and was run once under the same conditions as sparrow. For each instance, the best $\mathbb{E}(\rho)$ is highlighted in bold.
Figure 9: Distribution of sparrow’s solution quality $\rho$ across the academic benchmark instances. Each dot represents an individual solution from the experiments composing Table 2. The expected solution qualities of ROMA - Sato et al. (2019), GCS - Elkeran (2013), FLD - Wang et al. (2017) and ELS - Leung et al. (2012) algorithms are shown. A star marks the previous singular best-known solution for each instance. JAKOBS1 is omitted since it virtually always produced the same density.
@@
Table 3: The best solution densities produced by sparrow for instances in the benchmark set, accompanied by the previous best-known density and the algorithm that produced it: l - Leung et al. (2012). e - Elkeran (2013), w - Wang et al. (2017), s - Sato et al. (2019). Best densities are highlighted in bold. Time limit is 20 minutes.
@@ 12 Real-world benchmarks & future research directions
We introduce ten new instances that span a variety of application domains and contain both homogeneous and heterogeneous sets of items, as well as narrow and wider strips. Two variants are provided for each instance: one which only allows discrete rotations in 90-degree increments and a second which allows continuous rotation. The full dataset and experimental results are available in sparrow’s repository.
@@
Figure 12 presents this new suite instances for the 2DISPP. Instances GARDEYN0 - 3 have been derived from Lallier et al. (2022), a large dataset of nesting jobs originating from fashion industry clients of Lectra [^4] https://lectra.com/en. Instances GARDEYN4 - 8 are based on shapes from metalworking and shipbuilding applications, provided by AlmaCAM [^5] https://almacam.com/. Finally, GARDEYN9 represents a direct-to-film printing context, contributed by Tetrinest [^6] https://tetrinest.com/.
@@ 13 Conclusions and reflections
Reproducibility is a cornerstone of the scientific method. Without it, researchers are left in an unhealthy and inefficient environment: forced to reinvent the wheel without the means to increment on prior work or even validate its results. We strongly believe that an academic paper coupled with a high-quality and open-source implementation delivers value significantly exceeding the sum of its parts. Thankfully, the ethos we are advocating is not swimming against the current. The field of Operations Research is undergoing a gradual cultural shift toward open-source practices (Petropoulos et al., 2024). Vidal (2022), for example, recently published an open-source implementation of their state-of-the-art algorithm for the capacitated vehicle routing problem. Many of the concerns we raise here are in fact echoing sentiments already expressed in Vidal (2022).

### v2

@@ 8.3 Balancing exploration and exploitation
1: $s_{i}\leftarrow$ current solution ; $e^{*}\leftarrow\infty$
2: for N_WORKERS times do $\triangleright$ parallelized
3:   move_items $\triangleright$ Alg. 5
@@
5:   if $e<e^{*}$ then
6:    $s^{*}\leftarrow$ current solution ; $e^{*}\leftarrow e$
7:   restore to $s_{i}$
8: restore to $s^{*}$
@@ 9 Solving the 2D irregular strip packing problem
1: $s^{*}\leftarrow$ explore $\triangleright$ Alg. 12
2: $s^{*}\leftarrow$ compress($s^{*}$) $\triangleright$ Alg. 13
3: return $s^{*}$
@@ 9.1 Exploration phase
1: $s^{*}\leftarrow$ construct initial solution and shrink strip by $\texttt{R}_{x}$
2: while $\texttt{TL}_{x}$ not reached do
3:   $s\leftarrow\texttt{separate}(\texttt{K}_{x},\texttt{N}_{x})$ $\triangleright$ Alg. 9
4:   if $s$ is feasible then
5:    $s^{*}\leftarrow s$
6:    shrink strip by $\texttt{R}_{x}$
7:    $\mathcal{S}\leftarrow\emptyset$
8:   else
9:    $\mathcal{S}\leftarrow\mathcal{S}\cup\{s\}$ $\triangleright$ pool of infeasible solutions
10:    $\hat{s}\leftarrow$ select from $\mathcal{S}$
11:    $s^{\prime}\leftarrow$ disrupt $\hat{s}$ by swapping two large items
12:    restore to $s^{\prime}$
13: return $s^{*}$
@@
The separation procedure (Algorithm 9) attempts to convert this infeasible solution into a feasible one. If successful, the returned solution $s$ replaces the incumbent solution $s^{*}$, after which the strip is once again shrunk. If unsuccessful, $s$ represents an infeasible local optimum that is added to a pool of solutions $\mathcal{S}$. For the next separation attempt, a random solution $\hat{s}$ is chosen from $\mathcal{S}$, with those closer to feasibility (Equation 8) having a greater chance of being selected. A new starting solution $s^{\prime}$ is then created by swapping two large items from $\hat{s}$. This ensures $s^{\prime}$ inherits many high-quality placements from $\hat{s}$ while, at the same time, being sufficiently disrupted so it can escape from the local optimum. This procedure continues until time limit $\texttt{TL}_{x}$ is reached, at which point the exploration phase returns the best feasible solution found $s^{*}$.
@@ 9.2 Compression phase
1: while $\texttt{TL}_{c}$ not reached do
2:   $\tau\leftarrow\text{elapsed time in phase}$
3:   $r\leftarrow\texttt{R}_{c}^{s}+(\texttt{R}_{c}^{e}-\texttt{R}_{c}^{s})\cdot(\tau/\texttt{TL}_{c})$
4:   restore to $s^{*}$ and shrink strip by $r$
5:   $s^{\prime}\leftarrow\texttt{separate}(\texttt{K}_{c},\texttt{N}_{c})$ $\triangleright$ Alg. 9
6:   if $s^{\prime}$ is feasible then
7:    $s^{*}\leftarrow s^{\prime}$
8: return $s^{*}$
@@
Shrink ratio $r$ is much smaller than during the exploration phase and decays linearly from $\texttt{R}_{c}^{s}$ to $\texttt{R}_{c}^{e}$ based on the phase’s remaining time limit $\texttt{TL}_{c}$. This means that the separate procedure only needs to resolve (progressively) less severe collisions, which it is often able to do without making major changes to the configuration of items. Another important difference with the exploration phase is that, before each new separation attempt, the best feasible solution $s^{*}$ is always restored and the strip is shrunk again. This causes the search process’ freedom to be significantly more restricted than during the exploration phase. The compression phase is squarely focused on finding out how much further the strip can be shrunk without deviating too far from the configuration of items provided by the exploration phase.
@@ 10 Implementation
The implementation of our 2D nesting algorithm is written in Rust and called sparrow. It is publicly available at: https://github.com/JeroenGar/sparrow. The codebase closely follows the structure of this paper, making it straightforward to understand how sparrow relates to the algorithms presented here. Throughout this project, we have strived for maximum computational performance, robustness and leanness of both the algorithm and its implementation.
@@
So far, this limitation concerning exact fits only surfaces in artificial academic instances featuring simple shapes and vertices with integer coordinates. Conditions that, generally, do not occur in real-world applications.
@@ 11.1 Experimental setup
We used a combination of manual tuning and hyperparameter optimization (bergstra2013making) to identify a suitable set of parameters for the algorithm. These parameters were optimized for the maximum expected performance across all instances from the academic benchmark set when subject to a 20-minute time limit. This single set of parameters was used for every instance in our experiments.
@@ 11.2 Performance analysis
To evaluate the performance of sparrow we will use academic benchmark instances from the ESICUP website [^3] https://www.euro-online.org/websites/esicup/data-sets. We refer to Table 3 in sato2019raster for an overview of these instances and their properties.
To provide a sense of the efficiency of the implementation and scale of operations, when using the aforementioned experimental setup, sparrow is able to evaluate samples (Algorithm 7) at a rate of approximately 2.4 million per second with SWIM, which represents the most complex instance in the academic benchmark set.
@@
The expected quality $\mathbb{E}(\rho)$ and interquartile range of sparrow’s solutions in comparison to the academic SotA algorithms for the 2DISPP and one open-source implementation are presented in Table 2. The distributions of the solutions from Table 2 are visualized as violin plots in Figure 9. One important thing to note is that the width of the strip for some of the most simple instances – SHAPES0, SHAPES1, SHAPES2, FU, JAKOBS1 and JAKOBS2 – was inflated by 0.01%. This minuscule inflation should not affect which configurations of items are possible, but suffices to avoid the issue concerning exact fits (described in Section 10), which would otherwise occur in these artificial instances. For academic purposes, sparrow’s best solution ever produced for each instance, alongside the previous best-known solution, is presented in Table 3 and shown in Figure 10. It is important to note that these ‘best’ solutions, although conceived under the same experimental conditions and time limit, are not necessarily elements of the sets of 100 independent runs from which the expected densities in Table 2 have been derived.
@@
Table 2: The expected solution quality $\mathbb{E}(\rho)$ on academic benchmarks obtained by sparrow, four algorithms from the literature: ROMA - sato2019raster, GCS - elkeran2013new, FLD - wang2017flexible, ELS - leung2012extended and one open-source implementation: PS - fontan. For sparrow, quartiles (Q1, Q3) and the margin of error ($\pm$) on $\mathbb{E}$, defined as the half-width of the 95% confidence interval computed via non-parametric bootstrapping, are reported. The reported values for sparrow result from 100 independent runs, each with a time limit of 20 minutes. PS is a deterministic algorithm and was run once under the same conditions as sparrow. For each instance, the best $\mathbb{E}(\rho)$ is highlighted in bold.
@@
Table 3: The best solution densities ever produced by sparrow for instances in the benchmark set, accompanied by the previous best-known density and the algorithm that produced it: l - leung2012extended. e - elkeran2013new, w - wang2017flexible, s - sato2019raster. Best densities are highlighted in bold. Time limit is 20 minutes.
Figure 9: Distribution of sparrow’s solution quality $\rho$ across the academic benchmark instances. Each dot represents an individual solution from the experiments composing Table 2. The expected solution qualities of ROMA - sato2019raster, GCS - elkeran2013new, FLD - wang2017flexible and ELS - leung2012extended algorithms are shown. A star marks the previous singular best-known solution for each instance. JAKOBS1 is omitted since it virtually always produced the same density.
@@ 12 Real-world benchmarks & future research directions
We introduce ten new instances that span a variety of application domains and contain both homogeneous and heterogeneous sets of items, as well as narrow and wider strips. Every item is represented as a simple polygon and two variants are provided for each instance: one which only allows discrete rotations in 90-degree increments and a second which allows continuous rotation. The full dataset and experimental results are available in sparrow’s repository and in the ESICUP datasets  [^4] https://github.com/ESICUP/datasets.
@@
Figure 12 presents this new suite instances for the 2DISPP. Instances GARDEYN0 - 3 have been derived from corentin_lallier_2022_7030786, a large dataset of nesting jobs originating from fashion industry clients of Lectra [^5] https://lectra.com/en. Instances GARDEYN4 - 8 are based on shapes from metalworking and shipbuilding applications, provided by AlmaCAM [^6] https://almacam.com/. Finally, GARDEYN9 represents a direct-to-film printing context, contributed by Tetrinest [^7] https://tetrinest.com/.
@@ 13 Conclusions and reflections
Reproducibility is a cornerstone of the scientific method. Without it, researchers are left in an unhealthy and inefficient environment: forced to reinvent the wheel without the means to increment on prior work or even validate its results. We strongly believe that an academic paper coupled with a high-quality and open-source implementation delivers value significantly exceeding the sum of its parts. Thankfully, the ethos we are advocating is not swimming against the current. The field of Operations Research is undergoing a gradual cultural shift toward open-source practices (petropoulos2024operational). vidal2022hybrid, for example, recently published an open-source implementation of their state-of-the-art algorithm for the capacitated vehicle routing problem. Many of the concerns we raise here are in fact echoing sentiments already expressed in vidal2022hybrid.